The following data pertain to the Hercules Tire and Rubber Company for the month of May. 60,000 75,000 10,000 75,000 $ 135,00Missing Values Equivalent Units Cost per EU Total Cost of End WIP Calculate the missing values from the table given above. (RMissing Values Equivalent Units Cost per EU Total Cost of End WIP Calculate the Equivalent Units. Physical Units Equivalent UMissing Values Equivalent Units Cost per EU Total Cost of End WIP Calculate the Cost per EU. (Round Cost per equivalent unitMissing Values Equivalent Units Cost per EU Total Cost of End WIP Calculate the total cost of ending work-in-process. (Round
Answer #1

1.

Work in Process, May 1 (units) 15000 =75000-60000
Units completed and transferred 65000 =75000-10000
Total Equivalent units : conversion 71000 =65000+10000*60%
Work in Process, May 1 conversion $                37,500 =172500-135000
Costs incurred during may : direct material $              570,000 =1402250-832250
Cost per equivalent unit : conversion $                   12.25
Cost of goods completed and transferred out $          1,407,250
Cost remaining in ending work in process inventory : direct material $                94,000

2.

Equivalent units
Physical units Direct Material Conversion
Units completed and transferred out 65000 65000 65000
Work in Process, May 31 10000 10000 6000
Total units accounted for 75000
Total Equivalent units 75000 71000

3.

Direct Material Conversion Total
Work in Process, May 1 $              135,000 $                  37,500 $             172,500
Costs incurred during May $              570,000 $                832,250 $          1,402,250
Total Costs to account for $              705,000 $                869,750 $          1,574,750
Equivalent units 75000 71000
Cost per Equivalent unit $                     9.40 $                    12.25 $                  21.65


4.

No of Equivalent units Cost per Equivalent units Total Cost
Cost of goods completed and transferred 65000 $                    21.65 $          1,407,250
Cost remaining in May 31 work in process
Direct Material 10000 $                       9.40 $                94,000
Conversion 6000 $                    12.25 $                73,500
Total Cost of May 31, work in process $             167,500